Assuming a frictionless pipeline, how many litres of water can be raised in a 2500m vertical shaft in 35 minutes by a 550kw pump? (1litre= 1kg)
Answer #1

Energy Produced by pump to raise the water is

E=Pt =(550*103)(35*60) =1.155*109 Joules

Since E=mgh

=>m =E/gh = 1.155*109/9.8*2500

m=4.714*104 litres